# Multitask Learning Can Improve Worst-Group Outcomes

### 05 Dec 2023 v1

@@ Abstract
In order to create machine learning systems that serve a variety of users well, it is vital to not only achieve high average performance but also ensure equitable outcomes across diverse groups. However, most machine learning methods are designed to improve a model’s average performance on a chosen end task without consideration for their impact on worst group error. Multitask learning (MTL) is one such widely used technique. In this paper, we seek not only to understand the impact of MTL on worst-group accuracy but also to explore its potential as a tool to address the challenge of group-wise fairness. We primarily consider the common setting of fine-tuning a pre-trained model, where, following recent work (Gururangan et al., 2020; Dery et al., 2023), we multitask the end task with the pre-training objective constructed from the end task data itself. In settings with few or no group annotations, we find that multitasking often, but not always, achieves better worst-group accuracy than Just-Train-Twice (JTT; Liu et al. (2021)) – a representative distributionally robust optimization (DRO) method. Leveraging insights from synthetic data experiments, we propose to modify standard MTL by regularizing the joint multitask representation space. We run a large number of fine-tuning experiments across computer vision and natural language and find that our regularized MTL approach consistently outperforms JTT on both worst and average group outcomes. Our official code can be found here: https://github.com/atharvajk98/MTL-group-robustness.
@@ 1 Introduction
As machine learning systems exert ever-increasing influence on the real world, it is paramount that they not only perform well on aggregate but also exhibit equitable outcomes across diverse subgroups characterized by attributes like race (Buolamwini & Gebru, 2018; Liang et al., 2021), gender (Buolamwini & Gebru, 2018; Srinivasan & Bisk, 2022) and geographic location (Jurgens et al., 2017; De Vries et al., 2019; Ayush et al., 2021). It is therefore important to understand the impact of widely-used machine learning techniques with respect to these desiderata. Multitask learning (MTL) (Caruana, 1997; Baxter, 2000; Ruder et al., 2019; Dery et al., 2021a) is one example of such a technique that features prominently in machine learning practitioners’ toolbox for improving a model’s aggregate performance. However, the effect of multitask learning on worst group outcomes is underexplored. In this paper, we both study the impact of MTL, as is, on worst group error, and also consider whether modifications can be made to improve its effect on worst group outcomes.
Traditionally, the problem of worst-group generalization has been tackled explicitly via methods such as distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Duchi & Namkoong, 2018; Hashimoto et al., 2018; Sagawa et al., 2020a). In contrast to traditional empirical risk minimization, DRO aims to minimize the worst-case risk over a predefined set of distributions (the uncertainty set). Defining the uncertainty set usually (but not always) requires access to group annotations. Since average performance based approaches like MTL are typically designed without consideration of group annotations, our focus will be on settings with limited-to-no group annotations. In these settings, there exist a number of Generalized Reweighting (GRW) algorithms for distributional robustness (Nam et al., 2020; Liu et al., 2021; Zhang et al., 2022; Nam et al., 2022; Qiu et al., 2023; Zhai et al., 2023; Izmailov et al., 2022). These approaches minimize the weighted average risk based on the weight assigned to each example. One such widely used method is the Just-Train-Twice (JTT) algorithm (Liu et al., 2021) which performs two model training runs: one to identify poorly performing examples and another run that upweights these examples. For our empirical explorations, we take JTT as a representative DRO method and use it to provide reference performance to situate our study of multitask learning.
@@
We focus our investigations of multitasking on the ubiquitous setting of fine-tuning a pre-trained model. Here, a common way to improve end task average performance is to multitask the end task with the pre-training objective constructed over the task data itself (Gururangan et al., 2020; Dery et al., 2021b). We intuit that this multitasking approach could improve robustness to worst group outcomes since previous work like Hendrycks et al. (2019; 2020) has established a favorable connection between pre-training and robustness (both adversarial and out-of-distribution) whilst other work by Mao et al. (2020) has demonstrated multitasking’s positive impact on adversarial robustness. We test our intuition by conducting preliminary experiments across a pair of computer vision and natural language tasks in two settings: one with limited group annotations and the other with none. Initial results (Table 1) reveal that multitasking shows promise, in that it can improve worst group outcomes over ERM and JTT, but these improvements are not consistent. We are therefore spurred to consider modifications to make it a more competitive tool against worst group outcomes.
@@
In order to build intuition about how to adapt MTL to target worst group error, we conduct controlled experiments on two-layer linear models trained from scratch on synthetic data. We borrow the synthetic data setup introduced by Sagawa et al. (2020b) in which training data consists of two majority groups where spurious features (features that are not required to robustly solve the end task) are predictive of the end task output, and two minority group where spurious features are uncorrelated with the output. Sagawa et al. (2020b) demonstrated that under certain conditions on the generative distribution of the input features, linear models trained on such data will provably rely on spurious features and thus suffer poor worst group error. Working with this simplified setup where poor group outcomes are easily inducible allows us to more incisively study multitasking’s effects. To perform multitasking in the synthetic setup, we instantiate reconstruction from noised input as our auxiliary task. This choice is partially***We will delve deeper into other motivations in Section 2 informed by the fact that many pre-training objectives like masked language modeling (MLM) (Devlin et al., 2018) and masked image auto-encoders (He et al., 2022; Tong et al., 2022) are based on input reconstruction. When training solely on this auxiliary task, we uncover that regularizing the pre-output layer of the model is critical for ensuring that the model upweights the core features (features that are required to robustly solve the end task) over the spurious ones. This leads us to the following recipe for improving worst-group error: regularized multitasking of the end task with the (appropriately chosen) auxiliary objective.
@@
Through a battery of experiments across natural language processing (NLP) and computer vision (CV) datasets, we demonstrate that multitasking the end task with the pre-training objective along with $\ell_{1}$ regularization on the shared, pre-prediction layer activations is competitive when pitted against state-of-the-art DRO approaches like Just-Train-Twice (Liu et al., 2021) and Bitrate-Constrained DRO (Setlur et al., 2023). Specifically, in settings where only validation group annotations are available, regularized MTL outperforms JTT and BR-DRO on $/3$ 3 and $2/3$ datasets, respectively. Our approach improves worst-group accuracy over ERM (by as much as $\sim 4\%$) and JTT (by $\sim 1\%$) in settings where group annotations are completely unavailable. Moreover, regularized MTL consistently outperforms both ERM and JTT on average performance, regardless of whether group annotations are available or not. Therefore, within the prevailing framework of utilizing pre-trained models for downstream fine-tuning, our results demonstrate that regularized multitask learning can be a simple yet versatile and robust tool for improving both average and worst-group outcomes.
@@ 2 Informal motivation for our regularized MTL method
Why would we expect multitask learning to help mitigate worst group outcomes? It would be naive to assume that multitasking the end task with any auxiliary task would prevent poor group outcomes. In order to better understand intuitively which auxiliary tasks may be helpful, we first provide an example using the data generation process and linear model setup presented in Sagawa et al. (2020b)’s previous work on the effect of spurious and core features on worst-group accuracy.
@@
Sagawa et al. (2020b) describe a simple data-generating distribution that defines, for each example, a label $y\in\{-1,1\}$, a spurious attribute $s\in\{-1,1\}$, and features $x$. The features are described as either core features $x_{\text{core}}$ if they are associated with the label $y$, or spurious features $x_{\text{spur}}$ if they are associated with the spurious attribute $a$:
@@
We can then define a linear model, parameterized by $\hat{\mathbf{w}}$, that predicts the label given the features
@@
The cross-product of the space of possible labels $y=\pm 1$ and spurious attributes $s=\pm 1$ divides samples generated from this distribution into four groups. When some of the groups are more frequent than others in the training data, a correlation between $\{y,s\}$ is created. Further still, in the presence of the above correlation, if the spurious features have lower variance with respect to the data generating process (Equation 1) i.e $\sigma^{2}_{\mathrm{spur}}\leq\sigma^{2}_{\mathrm{core}}$, linear models will tend to rely more on (assign higher weight to) the spurious features over the core ones (Sagawa et al., 2020b). This learned reliance on spurious features – instead of core features that are truly predictive of the label – results in poor worst-group error.
@@
For instance, in the example above, if we add noise with a constant variance of $\sigma^{2}_{\mathrm{noise}}$ over each dimension, it results in noised inputs that have variances $\tilde{\sigma}^{2}_{\mathrm{spur}}=\left(\sigma^{2}_{\mathrm{spur}}+\sigma^{2}_{\mathrm{noise}}\right)\leq\tilde{\sigma}^{2}_{\mathrm{core}}=\left(\sigma^{2}_{\mathrm{core}}+\sigma^{2}_{\mathrm{noise}}\right)$ per spurious and core feature dimension respectively. Under the assumption that both true labels $y=1$ and $y=-1$ are equally probable, and in the simplest case where we are reconstructing features independently of each other with a linear predictor $\hat{x}_{i}=\bm{w}_{i}\tilde{x}_{i}$ (where $\tilde{x}$ is the noised input), the Bayes optimal weight on a feature $i$ would be (see Appendix A.2 for the full proof):
@@
Even given an auxiliary task with the above property of preferring core to spurious features (under $\sigma^{2}_{\mathrm{spur}}\leq\sigma^{2}_{\mathrm{core}}$), a model with sufficient capacity can still rely on spurious features for solving the end task. We can incentivise the model to mostly use core features by applying sufficient regularization to the parts of the model that are shared between the two tasks (such as shared feature extractors). The restricted capacity encourages the model to rely on features that would cause it to do well on both tasks, which would be the core features.
Based on the intuition established in this section, we propose a simple yet effective method for improving worst-group outcomes: multitasking the end task with the pre-training objective – which tend to be reconstruction tasks – whilst regularizing the shared (pre-prediction) layer. In Section 3 we will test this intuition though synthetic data experiments and in Sections 4 and 5, demonstrate its empirical efficacy through natural data experiments.
@@ 3 Synthetic Data Experiments
Figure 1: Visualization of synthetic training data (1,000 points). We generate training data according to the generative process specified in Equation 1 and the parameters specified in Section 3.2
We begin with an empirical study of a simplified setting of training a two-layer linear model on synthetic data. By a exploring such a setting, we hope ground our informal intuition from Section 2 in experiment.
@@ 3.1 Data Generating Distribution
As an instantiation, we consider end task data defined by $\mathbf{T}_{\mathrm{end}}=\{(x_{i},y_{i})\}_{i\in N}$ where we have $N$ total samples. Here $d_{c}=1;\leavevmode\nobreak\ d_{s}=1\implies d=2$. The data is dominated by samples where $\{s=y\}$ and thus we have two majority groups $\mathbf{G}_{s=y=1},\mathbf{G}_{s=y=-1}$ each with $\frac{n_{\mathrm{maj}}}{2}$ samples. The two minority groups are when $\{s=-y\}$ each with $\frac{n_{\mathrm{min}}}{2}$: $\mathbf{G}_{s=-y=1},\mathbf{G}_{s=-y=-1}$ Due to the fact that $n_{\mathrm{maj}}>n_{\mathrm{min}}$, the attribute $s$ is highly correlated with the label $y$ in the training data and is thus a spurious feature when considering the true data generation distribution. The end task is to predict the true label $y_{i}$ from the given input data $x_{i}$.
@@ 3.2 Training on the end task only
Figure 1 shows data sampled from the generative process described in Equation 1. We produce 1,000 samples in $\mathbf{R}^{2}$ with $\sigma^{2}_{\mathrm{core}}=0.6$ and $\sigma^{2}_{\mathrm{spur}}=0.1$. $n_{\mathrm{min}}=100$ and $n_{\mathrm{max}}=900$, making the spurious feature highly correlated with the true label.
@@
We train on $\mathbf{T}_{\mathrm{end}}$ only to confirm that the resulting model has poor worst group outcomes. Since we will eventually be performing multitasking, we use a two layer linear model where the first layer is a linear featurizer – that will eventually be shared between all tasks being multitasked – and the second layer is a prediction head dedicated to the end task. This shared featurizer but separate head architecture is common in modern multitask learning (Yu et al., 2020; Michel et al., 2021; Dery et al., 2021a).
@@
Note that we have effectively parameterized a linear model with decomposed formulation which will be helpful once we proceed to multitasking. Let $\sigma$ be the sigmoid function. The end task loss is binary cross-entropy with $\ell_{2}$ regularization on $w^{\mathrm{end}}$:
$\mathcal{L_{\mathrm{end}}}\left(w^{\mathrm{end}},a\right)=\frac{1}{N}\sum_{(x_{i},y_{i})\in\mathbf{T}_{\mathrm{end}}}\bigg{[}y_{i}\cdot\log\left(\sigma\left(y^{\mathrm{end}}_{\mathrm{pred}}\right)\right)+(1-y_{i})\cdot\log\left(1-\sigma\left(y^{\mathrm{end}}_{\mathrm{pred}}\right)\right)\bigg{]}+\frac{\lambda}{2}\|w^{\mathrm{end}}\|^{2}$
We fit the model solely to the end task by running batched stochastic gradient descent on $\mathcal{L_{\mathrm{end}}}$. We use a batch size of 64, learning rate of $10^{-3}$ and run for 500 epochs. $\lambda=1$ and we use 100 generated points as validation data for model selection. As can be seen in Figure 2, training on the end task only can result in a predictor that achieves poor worst group error. This occurs even with varying the norm of the featurizer parameter $\mathbf{a}$.

### 28 Feb 2024 v2

@@ Abstract
In order to create machine learning systems that serve a variety of users well, it is vital to not only achieve high average performance but also ensure equitable outcomes across diverse groups. However, most machine learning methods are designed to improve a model’s average performance on a chosen end task without consideration for their impact on worst group error. Multitask learning (MTL) is one such widely used technique. In this paper, we seek not only to understand the impact of MTL on worst-group accuracy but also to explore its potential as a tool to address the challenge of group-wise fairness. We primarily consider the standard setting of fine-tuning a pre-trained model, where, following recent work (Gururangan et al., 2020; Dery et al., 2023), we multitask the end task with the pre-training objective constructed from the end task data itself. In settings with few or no group annotations, we find that multitasking often, but not consistently, achieves better worst-group accuracy than Just-Train-Twice (JTT; Liu et al. (2021)) – a representative distributionally robust optimization (DRO) method. Leveraging insights from synthetic data experiments, we propose to modify standard MTL by regularizing the joint multitask representation space. We run a large number of fine-tuning experiments across computer vision and natural language processing datasets and find that our regularized MTL approach consistently outperforms JTT on both average and worst-group outcomes. Our official code can be found here: https://github.com/atharvajk98/MTL-group-robustness.
@@ 1 Introduction
As machine learning systems exert ever-increasing influence on the real world, it is paramount that they not only perform well on aggregate but also exhibit equitable outcomes across diverse subgroups characterized by attributes like race (Buolamwini & Gebru, 2018; Liang et al., 2021), gender (Buolamwini & Gebru, 2018; Srinivasan & Bisk, 2022) and geographic location (Jurgens et al., 2017; De Vries et al., 2019; Ayush et al., 2021). Therefore, it is important to understand the impact of widely-used machine learning techniques with respect to these desiderata. Multitask learning (MTL) (Caruana, 1997; Baxter, 2000; Ruder et al., 2019; Dery et al., 2021a) is one example of such a technique that features prominently in machine learning practitioners’ toolbox for improving a model’s aggregate performance. However, the effect of multitask learning on worst group outcomes is underexplored. In this paper, we both study the impact of MTL, as is, on worst group error and also consider whether modifications can be made to improve its effect on worst group outcomes.
Traditionally, the problem of worst-group generalization has been tackled explicitly via methods such as distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Duchi & Namkoong, 2018; Hashimoto et al., 2018; Sagawa et al., 2020a). In contrast to traditional empirical risk minimization, DRO aims to minimize the worst-case risk over a predefined set of distributions (the uncertainty set). Defining the uncertainty set usually (but not always) requires access to group annotations. Since average performance based approaches like MTL are typically designed without consideration of group annotations, our focus will be on settings with limited-to-no group annotations. In these settings, there exist a number of generalized Reweighting (GRW) algorithms for distributional robustness (Nam et al., 2020; Liu et al., 2021; Zhang et al., 2022; Nam et al., 2022; Qiu et al., 2023; Zhai et al., 2023; Izmailov et al., 2022). These approaches minimize the weighted average risk based on the weight assigned to each example. One such widely used method is the Just-Train-Twice (JTT) algorithm (Liu et al., 2021), which performs two model training runs: one to identify poorly performing examples and another run that upweights these examples. For our empirical explorations, we take JTT as a representative DRO method and use it to provide reference performance to situate our study of multitask learning.
@@
We focus our investigations of multitask learning on the ubiquitous setting of fine-tuning a pre-trained model. Here, a common way to improve end task average performance is to multitask the end task with the pre-training objective constructed over the task data itself (Gururangan et al., 2020; Dery et al., 2021b). We intuit that this multitasking approach could improve robustness to worst group outcomes since previous work like Hendrycks et al. (2019; 2020); Mao et al. (2020) has established a favorable connection between pre-training and robustness (both adversarial and out-of-distribution). We test our intuition by conducting preliminary experiments across a pair of computer vision and natural language tasks in two settings: one with limited group annotations and the other with none. Initial results (Table 1) reveal that multitasking shows promise in that it can improve worst group outcomes over ERM and JTT/ However, these improvements are not consistent. Therefore, we are spurred to consider modifications to make it a more competitive tool against worst group outcomes.
@@
In order to build intuition about how to adapt MTL to target the worst group error, we conduct controlled experiments on two-layer linear models trained from scratch on synthetic data. We borrow the synthetic data setup introduced by Sagawa et al. (2020b) in which training data consists of two majority groups, where spurious features (features that are not required to robustly solve the end task) are predictive of the end task output, and two minority groups, where spurious features are uncorrelated with the output. Sagawa et al. (2020b) demonstrated that under certain conditions on the generative distribution of the input features, linear models trained on such data would probably rely on spurious features and thus, have poor worst group error. Working with this simplified setup where worst-group outcomes are easily inducible allows us to study MTL’s effects more incisively. We instantiate reconstruction from noised input as our auxiliary task to perform multitask learning in the synthetic setup. This choice is partially***We will delve deeper into other motivations in Section 2 informed by the fact that many pre-training objectives like masked language modeling (MLM) (Devlin et al., 2018) and masked image modeling (He et al., 2022; Tong et al., 2022) are based on input reconstruction. When training solely on this auxiliary task, we uncover that regularizing the pre-output layer of the model is critical for ensuring that the model upweights the core features (features required to robustly solve the end task) over the spurious ones. This leads us to the following recipe for improving worst-group error: regularized multitask learning of the end task with the (appropriately chosen) auxiliary objective.
@@
Through a battery of experiments across natural language processing (NLP) and computer vision (CV) datasets, we demonstrate that multitasking the end task with the pre-training objective along with $\ell_{1}$ regularization on the shared, pre-prediction layer activations is competitive when pitted against state-of-the-art DRO approaches like JTT (Liu et al., 2021) and Bitrate-Constrained DRO (Setlur et al., 2023). Specifically, in settings where only validation group annotations are available, regularized MTL outperforms JTT and BR-DRO on $/3$ 3 and $2/3$ datasets, respectively. Our approach improves worst-group accuracy over ERM (by as much as $\sim 4\%$) and JTT (by $\sim 1\%$) in settings where group annotations are completely unavailable. Moreover, regularized MTL consistently outperforms both ERM and JTT on average performance, regardless of whether group annotations are available or not. Thus, within the prevailing framework of utilizing pre-trained models for downstream fine-tuning, our results demonstrate that regularized multitask learning can be a simple yet versatile and robust tool for improving both average and worst-group outcomes.
@@ 2 Informal motivation for our regularized MTL method
Problem Setup / Preliminaries:
Let each input example $x\in X$ have a classification label $y\in Y$ and a demographic attribute $s\in S$. Each group $g=(s,y)\in G$ is defined by the label $y$ and the attribute $s$, such that $s$ spuriously correlates with $y$. Thus, the sample space of $G$ is the Cartesian product of $Y$ and $S$. Our goal is to learn a model that minimizes the worst-group error. We evaluate the models based on their worst group accuracy (WGA), i.e., the minimum predictive accuracies of our model across all groups. We are interested in the setting where the spurious attribute $s$, and consequently, the group identity $g$ are unavailable (or available to a very limited degree) at training time, as annotating spurious attributes is typically expensive.
Why would we expect multitask learning to help mitigate worst group outcomes?
It would be naive to assume that multitasking the end task with any auxiliary task would prevent poor group outcomes. In order to better understand intuitively which auxiliary tasks may be helpful, we first provide an example using the data generation process and linear model setup presented in Sagawa et al. (2020b)’s work on the effect of spurious and core features on worst-group accuracy.
@@
Sagawa et al. (2020b) describe a simple data-generating distribution that defines, for each example, a label $y\in\{-1,1\}$, a spurious attribute $s\in\{-1,1\}$, and features $x$. The features are described as either core features $x_{\text{core}}$ if they are associated with the label $y$, or spurious features $x_{\text{spur}}$ if they are associated with the spurious attribute $s$:
@@
We can then define a linear model, parameterized by $\hat{\mathbf{w}}$, that predicts the label given the features.
@@
The cross-product of the space of possible labels $y=\pm 1$ and spurious attributes $s=\pm 1$ divides the samples generated from this distribution into four groups. When some groups are more frequent than others in the training data, a correlation between $\{y,s\}$ is created. Further still, in the presence of the above correlation, if the spurious features have lower variance with respect to the data generating process (Equation 1), i.e., $\sigma^{2}_{\mathrm{spur}}\leq\sigma^{2}_{\mathrm{core}}$, linear models will tend to rely more on (assign a higher weight to) the spurious features over the core ones (Sagawa et al., 2020b). This learned reliance on the spurious features – instead of the core features that are truly predictive of the label – results in poor worst-group error.
@@
For instance, in the example above, if we add noise with a constant variance of $\sigma^{2}_{\mathrm{noise}}$ over each dimension, it results in noised inputs that have variances $\tilde{\sigma}^{2}_{\mathrm{spur}}=\left(\sigma^{2}_{\mathrm{spur}}+\sigma^{2}_{\mathrm{noise}}\right)\leq\tilde{\sigma}^{2}_{\mathrm{core}}=\left(\sigma^{2}_{\mathrm{core}}+\sigma^{2}_{\mathrm{noise}}\right)$ per spurious and core feature dimension, respectively. Under the assumption that both true labels $y=1$ and $y=-1$ are equally probable, and in the simplest case where we are reconstructing features independently of each other with a linear predictor $\hat{x}_{i}=\bm{w}_{i}\tilde{x}_{i}$ (where $\tilde{x}$ is the noised input), the Bayes optimal weight on a feature $i$ would be (see Appendix A.3 for the full proof):
@@
Given an auxiliary task with the above property of preferring core to spurious features (under $\sigma^{2}_{\mathrm{spur}}\leq\sigma^{2}_{\mathrm{core}}$), a model with sufficient capacity can still rely on spurious features for solving the end task. We can incentivize the model to mostly use core features by applying sufficient regularization to the parts of the model that are shared between the two tasks (such as shared feature extractors). The restricted capacity encourages the model to rely on features that would cause it to do well on both tasks, which would be the core features.
Based on the intuition established in this section, we propose a simple yet effective method for improving worst-group outcomes: Multitasking the end task with the pre-training objective – which tends to be reconstruction tasks – while regularizing the shared (pre-prediction) layer. In Section 3, we will test this intuition through synthetic data experiments, and in Sections 4 and 5, we demonstrate its empirical efficacy through natural data experiments.
@@ 3 Synthetic Data Experiments
We initiate our investigation with an empirical study in a simplified context, involving the training of a two-layer linear model on synthetic data. This exploration is designed to substantiate the informal intuition introduced in Section 2 through concrete empirical findings.
@@ 3.1 Data Generating Distribution
Figure 1: Visualization of synthetic training data ($1000$ points).
As an instantiation, we consider end the task data defined by $\mathbf{T}_{\mathrm{end}}=\{(x_{i},y_{i})\}_{i\in N}$ where we have $N$ total samples. Here $d_{c}=1;\leavevmode\nobreak\ d_{s}=1\implies d=2$. The data is dominated by samples where $\{s=y\}$ and thus, we have two majority groups $\mathbf{G}_{s=y=1}$ and $\mathbf{G}_{s=y=-1}$, each with $\frac{n_{\mathrm{maj}}}{2}$ samples. The two minority groups are when $\{s=-y\}$ each with $\frac{n_{\mathrm{min}}}{2}$: $\mathbf{G}_{s=-y=1},\mathbf{G}_{s=-y=-1}$ Due to the fact that $n_{\mathrm{maj}}>n_{\mathrm{min}}$, the attribute $s$ is highly correlated with the label $y$ in the training data and thus, is a spurious feature when considering the true data generation distribution. The end task is to predict the true label $y_{i}$ from the given input data $x_{i}$.
Figure 1 shows data sampled from the generative process described in Equation 1. We produce $1000$ samples in $\mathbf{R}^{2}$ with $\sigma^{2}_{\mathrm{core}}=0.6$ and $\sigma^{2}_{\mathrm{spur}}=0.1$. $n_{\mathrm{min}}=100$ and $n_{\mathrm{max}}=900$, making the spurious feature highly correlated with the true label.
@@ 3.2 Training on the end task only
We train on $\mathbf{T}_{\mathrm{end}}$ only to confirm that the resulting model has poor worst-group outcomes. Since we will eventually be do multitasking, we use a two layer linear model where the first layer is a linear featurizer – that will eventually be shared between all tasks being multitasked – and the second layer is a prediction head dedicated to the end task. This shared featurizer but separate head architecture is common in modern multitask learning (Yu et al., 2020; Michel et al., 2021; Dery et al., 2021a).
@@
Note that we have effectively parameterized a linear model with decomposed formulation which will be helpful once we proceed to multitasking. The end task loss is binary cross-entropy with $\ell_{2}$ regularization on $w^{\mathrm{end}}$ is given by the following equation where $\sigma$ be the sigmoid function:
$\mathcal{L_{\mathrm{end}}}\left(w^{\mathrm{end}},a\right)=\frac{1}{N}\sum_{(x_{i},y_{i})\in\mathbf{T}_{\mathrm{end}}}\bigg{[}y_{i}\cdot\log\left(\sigma\left(y^{\mathrm{end}}_{\mathrm{pred}}\right)\right)+(1-y_{i})\cdot\log\left(1-\sigma\left(y^{\mathrm{end}}_{\mathrm{pred}}\right)\right)\bigg{]}+\frac{\lambda}{2}\|w^{\mathrm{end}}\|^{2}$ (5)
We fit the model solely to the end task by running batched stochastic gradient descent on $\mathcal{L_{\mathrm{end}}}$. We use a batch size of $64$, learning rate of $10^{-3}$, $\lambda=1$, and $500$ epochs. We use $100$ generated points as validation data for model selection. As can be seen in Figure 2, training on the end task only can result in a predictor that achieves poor worst group error. This occurs even with varying the norm of the featurizer parameter $\mathbf{a}$.